# Deep learning methods for inverse problems using connections between proximal operators and Hamilton–Jacobi equations

### v1

@@ 5 Numerical results
We evaluate Learned Proximal Networks (LPNs) for approximating the proximal operators of nonconvex and concave priors. While LPNs [fang2024whats] are theoretically grounded in convex analysis (parameterizing the proximal operator as the gradient of a convex potential $\psi$), these experiments investigate their behavior when trained on data generated from fundamentally nonconvex and concave landscapes. All experiments utilize the official LPN implementation. The network is trained via supervised learning, minimizing the mean squared error (MSE) or L1 loss between the network output and the true value. We use an LPN with $2$ layers and $256$ hidden units using Softplus activation ($\beta=5$) to ensure $C^{2}$ smoothness. The model is trained using the Adam optimizer with a starting learning rate of $10^{-3}$ and decreased by a factor of $10^{-1}$ at every $10^{5}$ epochs for a total of $5\times 10^{5}$ epochs.
The data generation process for all experiments is as follows: $N$ samples ($y_{i}$) are drawn uniformly from the hypercube $a,a]^{d}$, where $a$ is chosen to be $4$ and $d$ is the dimension, equal $2,4,8,16,32$ and $64$. $N=3\times 10^{4}$ is chosen for $d=2,4$, $N=3\times 10^{4}$ is chosen for $d=8,16$, and $N=4\times 10^{4}$ is chosen for $d=32,64$.
We also trained a second LPN to recover the prior at arbitrary points and compare its performance to the “invert” method (find $y$ such that $f_{\theta}(y)=x$) used in [fang2024whats] for recovering the prior from its proximal. Our second LPN is based on the relationship that the non-convex prior $J(x)$ can be approximated using the convex conjugate of the learned potential $\psi(y)$. Specifically, we compute:
(27) $J(x)\approx G(x)-\frac{1}{2}\|x\|^{2}$
where $G(x)=\psi^{*}(x)$ represents the convex conjugate of the potential $\psi_{\theta}(y)$ learned by the first LPN. We generate a new dataset $\{(x_{k},G_{k})\}$ using the trained first LPN $\psi_{\theta}$: (i) The gradients of the first network evaluated at the original sample points $y_{i}$,
(28) $x_{k}=\nabla_{y}\psi_{\theta}(y_{i}),$
and (ii) the values of the Legendre transform corresponding to each point,
(29) $G_{k}=\langle x_{k},y_{i}\rangle-\psi_{\theta}(y_{i}).$
The network $\phi_{G}$ is trained to map the gradients $x_{k}$ to the conjugate values $G_{k}$ by minimizing the Mean Squared Error (MSE). The optimization is performed using the Adam optimizer with the same parameters as used in the first LPN. Once the second LPN is trained, the estimated non-convex prior $\hat{J}(x)$ is recovered via
(30) $\hat{J}(x)=\phi_{G}(x)-\frac{1}{2}\|x\|^{2}.$
5.1 Convex prior
We will benchmark our approach with the prior $J(\bm{x})=\left\|{\bm{x}}\right\|_{1}$. For this example, we have
$\operatorname*{arg\,min}_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+\left\|{\bm{y}}\right\|_{1}\right\}$ $=\cup_{j=1}^{n}\operatorname*{arg\,min}_{y_{j}\in\mathbb{R}}\left\{\frac{1}{2t}(x_{j}-y_{j})^{2}+|y_{j}|\right\}$
@@ 5.1 Convex prior
With this, we can evaluate $S(\bm{x},t)$ and the LPN function $\bm{x}\mapsto\Psi(\bm{x})\coloneqq\frac{1}{2}-tS(\bm{x},t)$.
Table 1: Mean square errors of LPN $\psi$ and prior $J$ with 2 layers and 256 neurons in the convex L1 prior example.
Dimension LPN ($\psi$) Prior ($J$)
@@
Figure 1: The cross sections of the convex function $\psi(x)$ for dimension 88 (top). The bottom row compares the cross sections of the prior function from “invert LPN” (left) and our trained second LPN method (right).
5.2 Non-convex prior
Minplus algebra example
For this example, the prior is
@@ 5.2 Non-convex prior
We use $\mu_{1}=(1,0,\dots,0)$, $\mu_{2}=\bm{1}/\sqrt{n}$, and $\sigma_{1}=\sigma_{2}=1.0$.
Table 2: Mean square errors of LPN $\psi$ and prior $J$ with 2 layers and 256 neurons in the min-plus example.
Dimension LPN ($\psi$) Prior ($J$)
@@
Figure 2: The cross sections of the convex function $\psi(x)$ for dimension 88 (top). The bottom row compares the cross sections of the prior function from “invert LPN” (left) and our trained second LPN method (right).
Figure 3: The cross sections of the convex function $\psi(x)$ for dimension $32$ (top). The bottom row compares the cross sections of the prior function from “invert LPN” (left) and our trained second LPN method (right)
@@ 5.3 Concave prior
For this example, we use
@@
This is actually challenging example because, technically, $J$ is not uniformly Lipschitz continuous. (Although numerically we can get around this by “Huberizing” the prior.) We use this prior because we have an exact solution for this problem. It’s also a bit challenging for a convex LPN network because according to the theory of HJ PDEs, the function $J+\frac{1}{4}\left\|{\bm{x}}\right\|_{2}^{2}$ is convex, and $J+\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}$ is strongly convex, so an LPN (which is not inherently S.C.) may not be able to detect the strong convexity and makes this function more challenging to learn.
Table 3: Mean square errors of LPN $\psi$ and prior $J$ with 2 layers and 256 neurons in the concave prior example.
Dimension LPN ($\psi$) Prior ($J$)
@@
Figure 4: The cross sections of the prior function for $8$ dimension from “invert LPN” (left) and our trained second LPN method (right)
@@ 5.4 Negative $\ell_{1}$ norm
See Equation Eq. 32 for the minimum value and Eq. 33 for the proximal value. For this example, we consider $J(\bm{y})=-\left\|{\bm{y}}\right\|_{1}$. Let $n=1$ for simplicity and consider the one-dimensional problem
(31) $S(x,t)=\min_{y\in\mathbb{R}}\left\{\frac{1}{2t}(x-y)^{2}-|y|\right\}.$
The function $y\mapsto(x-y)^{2}/2t-|y|$ is differentiable everywhere except at $y=0$. A stationary point of this function satisfies
[…]
If $x>t$, the only minimum is $x+t$, in which case we have
$S(x,t)=\frac{1}{2t}(x+t-x)^{2}-|x+t|=\frac{t}{2}-(x+t)=-\frac{t}{2}-x.$
If $0<x\leqslant t$, there are two local minimums, 0 and $x+t$, but the global minimum is attained at $x+t$, again yielding $S(x,t)=\frac{t}{2}-x$. If $x=0$, we have three local minimums: t-t, 0, and $t$. The global minimum is attained at either t-t or $t$, yielding $S(0,t)=-\frac{t}{2}$. If $t\leqslant x<0$, there are two local minimums, 0 and xtx-t, but the global minimum is attained at xtx-t, yielding $S(x,t)=-\frac{t}{2}+x$. If x<tx<-t, the only minimum is xtx-t, in which case we have $S(x,t)-\frac{t}{2}+x$. Hence
(32) $S(x,t)=-\frac{t}{2}-|x|.$
In particular, its gradient in $x$ is given by
$\nabla_{x}S(x,t)\in\begin{cases}1&\ \text{if $x<0$},\\ [-1,1]&\ \text{if $x=0$},\\ -1&\ \text{if $x>0$}.\end{cases}$
Moreover,
(33) $\operatorname*{arg\,min}_{y\in\mathbb{R}}\left\{\frac{1}{2t}(x-y)^{2}-|y|\right\}=\begin{cases}x-t&\ \text{if $x<0$},\\ [-t,t]&\ \text{if $x=0$},\\ x+t&\ \text{if $x>0$}.\end{cases}$
@@
Figure 5: The cross sections of the convex function $\psi(x)$ for dimension 44 (left) and $8$ (right).
The Tables 1, 2, 3, and 4 quantify the scalability of the LPN approach across dimensions ranging from 22D to 6464D. The results indicate that the method performs with high accuracy in lower dimensions (22D through 88D), achieving mean square errors in the range of $10^{-7}$ to $10^{-4}$. However, performance degrades slightly as the dimensionality increases, particularly at 1616D and 3232D, where the error spikes to a range of $10^{-3}$ and $10^{-1}$. The higher dimensions did not work too well, which might be because we did not train for long enough and also used a simple architecture (which ought to be more intricate for the higher dimensional problems). While the error for the recovered prior $J$ is generally slightly higher than that of LPN $\psi$, this is expected given the added complexity of recovering the non-convex, non-smooth, or concave functions. However, the errors generally remain low, validating our method’s effectiveness even in high-dimensional spaces.
The top rows of Figures 1,2, 3 and Figure 5 demonstrate that the LPN accurately learns the cross sections of the convex function $\psi(x)$ for dimension $4,8,$ and $32$, closely matching the reference function with the most significant variation in Figure 3 corresponding to dimension $32$. The bottom row of these figures compares the “invert LPN” (left) method and our trained second LPN method (right). It is clear that in all cases, our direct method recovers the original non-smooth prior, as indicated by the sharp V-shaped reconstructions in Figure 1, the non-convex prior Figures 2 and 3, and quadratic concave prior Figure 4 despite its challenging nature.
Figure 5 represents the cross-sections of the LPN value function against the ground truth for dimensions $4$ (left) and $8$ (right). In both instances, the LPN-approximation curves exhibit a tight fit to the analytical reference, capturing the characteristic V-shape more closely and the non-smooth geometry of the underlying function. The errors for $64$ dimensions in all the examples are consistently lower than expected. In contrast, it does not yield a good visual approximation of the cross sections. We are unable to explain the reason for this behaviour in 6464D cases yet; however, we think it is probably due to the way we sample the hypercube.
@@ 6 Discussion
In this work, we leveraged the theory of viscosity solutions of HJ PDEs to develop novel deep learning numerical methods to learn, from data, the underlying prior of the proximal operator Eq. 2 yielding $(\bm{x},t)\mapsto S(\bm{x},t)$ defined in Eq. 1. Our approach built on the existing connections between proximal operators and HJ PDEs, crucially the fact that $(\bm{x},t)\mapsto S(\bm{x},t)$ is obtained from the solution to an HJ PDE, and in particular on the theory for the inverse problem for HJ equations. As discussed in Section 3, the theory for the inverse problem for HJ equations show that while there may be infinitely many priors that can recover Eq. 1, there is a natural choice, obtained by reversing the time in the HJ PDE Eq. 14 and using the value of the proximal operator $(\bm{x},t)\mapsto S(\bm{x},t)$ as initial condition. The resulting backward viscosity solution yields a prior $J_{\text{BVS}}$ that can reconstruct the $(\bm{x},t)\mapsto S(\bm{x},t)$ and also that is semiconvex. We considered the case where only samples of the proximal operators and its values were available in Section Section 4, and used techniques from max-plus algebra to derive some characterizations and errors property of $J_{\text{BVS}}$ with respect to convex functions approximating it from above. Finally, in Section 5 we proposed to learn the prior $J_{\text{BVS}}$ by training a convex neural network, specifically a learned proximal network, on a function of the form $\bm{y}\mapsto\tilde{J}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ from data $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$ via Eq. 17. We presented several numerical results that demonstrate the efficiency of our proposed method in high dimensions.
@@
Appendix A Calculations
A.1 Formal calculation of Eq. 19
We have
$\inf_{\bm{y}\in\mathbb{R}^{n}}$ $\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J_{\text{PAM}}(\bm{y})\right\}=\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+\right$.
$\qquad\qquad\qquad\left.\max_{k\in\{1,\dots,K\}}\left\{J_{\text{BVS}}(\bm{y}_{k})+\frac{1}{2t}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}-\frac{1}{2t}\left\|{\bm{x}_{k}-\bm{y}}\right\|_{2}^{2}\right\}\right\}$
$=\max_{k\in\{1,\dots,K\}}\left\{J_{\text{BVS}}(\bm{y}_{k})+\frac{1}{2t}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}+\right$.
$\qquad\qquad\qquad\left.\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}-\frac{1}{2t}\left\|{\bm{x}_{k}-\bm{y}}\right\|_{2}^{2}\right\}\right\}$
[…]
A.2 Formal calculation of Eq. 21

### v2

@@ 5 Numerical results
In this section, we leverage the theory of viscosity solutions of HJ PDEs to learn proximal operators of certain (non)convex priors. We used Learned Proximal Networks (LPNs) [fang2024whats] for this, which are designed to learn the proximal operator as the gradient of a convex potential. These experiments investigate their behavior when trained on data generated from fundamentally nonconvex and concave landscapes. All experiments used the official LPN implementation [fang2024whats].
We used LPNs with $2$ layers and $256$ hidden units using Softplus activation ($\beta=5$) to ensure $C^{2}$ smoothness. The models were trained via the Adam optimizer with a starting learning rate of $10^{-3}$ and decreased by a factor of $10^{-1}$ at every $10^{5}$ epochs for a total of $5\times 10^{5}$ epochs. We generated the data for all experiments as follow: $N$ samples $\{\bm{x}_{k}\}_{k=1}^{N}$ were drawn uniformly from the hypercube $a,a]^{d}$, where $a=4$ and $d=2,4,8,16,32$ and $64$. We chose $N=3\times 10^{4}$ $d=2$ and $4$, $N=3\times 10^{4}$ for $d=8$ and $16$, and $N=4\times 10^{4}$ for $d=32$ and $64$. In each experiment, we evaluated the minimum value of the proximal operator $S(\bm{x}_{k},1)$ at every sample.
The first phase of the experiment focused on learning the value function $S(\bm{x},1)$ associated with the HJ equation Eq. 3 with initial data $J$ at $t=1$. We encoded it into a convex potential function $\psi_{\text{NN}}(\bm{y};\theta)$, which we trained by minimizing the MSE loss
$\min_{\theta}\frac{1}{K}\sum_{k=1}^{K}\left(\psi_{\text{NN}}(\bm{x}_{k};\theta)-\left(\frac{1}{2}\left\|{\bm{x}_{k}}\right\|_{2}^{2}-S(\bm{x}_{k},1)\right)\right)^{2}.$
After training, we evaluated the trained value function using $\bm{x}\mapsto\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-\psi_{\text{NN}}(\bm{x};\theta)$.
@@
$\bm{y}_{k}=\nabla_{\bm{x}}\psi_{\text{NN}}(\bm{x}_{k};\theta),\quad J_{\text{BVS}}(\bm{y}_{k})\approx G_{k}\coloneqq\langle\bm{x}_{x},\bm{y}_{k}\rangle-\psi_{\text{NN}}(\bm{x}_{k},1)-\frac{1}{2}\left\|{\bm{y}_{k}}\right\|_{2}^{2}.$
We then trained the second LPN by minimizing the MSE loss
$\min_{\tilde{\theta}}\frac{1}{K}\sum_{k=1}^{K}\left(J_{\text{NN}}(\bm{y}_{k};\tilde{\theta})-G_{k}\right)^{2}.$
[…]
5.1 Convex prior: the $\ell_{1}$ norm
We begin with the convex prior $J(\bm{x})=\left\|{\bm{x}}\right\|_{1}$. For this example, the value function $S(\bm{x},t)$ can be evaluated using the soft-thresholding operator Eq. 17. The errors with respect the exact solutions of the first and second LPN, are presented in the Table 1. Some plots of cross sections of $\psi_{\text{NN}}$ with respect to the true convex function $\bm{x}\mapsto\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-S(\bm{x},1)$ are presented in Fig. 1 below. The results demonstrate that the first LPN learns fairly well the forward solution $\bm{x}\mapsto S(\bm{x},t)$ from samples of $\bm{x}$ and $S(\bm{x},1)$ even in high dimensions, though the MSE taper off as the dimension $d$ increases.
We wish to note that in dimension 64, the MSE for $\psi_{\text{NN}}$ reduced significantly compared to the previous dimensions. We are unsure as to why this is the case. It is possible that using a deeper neural network or incorporating more nodes may improve performance. The second LPN learns the true prior $\bm{x}\mapsto\left\|{\bm{x}}\right\|_{1}$ also fairly well, especially since it learning from the first, approximate LPN $\psi_{\text{NN}}$, without recourse to samples of the gradient $\bm{x}\mapsto\nabla_{\bm{x}}S(\bm{x},t)$.
Table 1: Mean square errors of LPN $\psi_{\text{NN}}$ and prior $J_{\text{NN}}$ with 2 layers and 256 neurons in the convex L1 prior example.
Dimension LPN ($\psi_{\text{NN}}$) Prior ($J_{\text{NN}}$)
@@ 5.1 Convex prior: the $\ell_{1}$ norm
Figure 1: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension 88 (left). The cross sections of the prior function learned from our trained second LPN method (right).
5.2 Non-convex priors
A min-plus algebra prior
For this example, we use a min-plus prior:
@@ 5.2 Non-convex priors
We used $\mu_{1}=(1,0,\dots,0)$, $\mu_{2}=\bm{1}/\sqrt{n}$, and $\sigma_{1}=\sigma_{2}=1.0$. We chose this prior despite it not being uniformly Lipschitz continuous, an assumption needed for having a unique viscosity solution to the HJ PDE, to check if an LPN could learn the prior regardless. We obtained the following results, presented in Table 2 and Fig. 2 and  3. The results are comparable to those obtained for the $\ell_{1}$ prior in Section 5.1. The MSE for the $d=64$ case also decreases dramatically, as in the $\ell_{1}$-convex prior, and we are unsure why this is the case. In any case, the second LPN performed reasonably well up in high dimensions.
Table 2: Mean square errors of LPN $\psi$ and prior $J_{\text{NN}}$ with 2 layers and 256 neurons in the min-plus example.
Dimension LPN ($\psi_{\text{NN}}$) Prior ($J_{\text{NN}}$)
@@
Figure 2: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension 88 (left). The cross sections of the prior function learned from our trained second LPN method (right).
Figure 3: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension $32$ (left). The cross sections of the prior function learned from our trained second LPN method (right).
@@ 5.3 Concave prior
For this example, we use the concave prior
@@
This is a challenging example because, as the min-plus prior, it is not uniformly Lipschitz continuous, although numerically since we constrain the domain to $4,4]^{d}$, we can get around this by “Huberizing” the prior outside $4,4]^{d}$. For this prior, we have $S(\bm{x},t)=-\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}$. After training, we obtained the results presented in Table 3 and Figs. 4. The results are similar to those obtained for the two previous prior, though we found this prior slightly more challenging to train. Again, for dimension $d=64$, the MSE for the first LPN decreased by a large factor.
Table 3: Mean square errors of LPN $\psi_{\text{NN}}$ and prior $J_{\text{NN}}$ with 2 layers and 256 neurons in the concave prior example.
Dimension LPN ($\psi_{\text{NN}}$) Prior ($J_{\text{NN}}$)
@@
Figure 4: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension 88 (left). The cross sections of the prior function learned from our trained second LPN method (right).
@@ 5.4 Negative $\ell_{1}$ norm
We finish with the nonconvex prior $J(\bm{x})=-\left\|{\bm{x}}\right\|_{1}$. As per examples 2.5 with $n=1$, we have
$S(\bm{x},t)=-\frac{n}{2}-\left\|{\bm{x}}\right\|_{1}.$
We present the results after training in Table 4 and some plots in Fig. 5. This example turned out to be very challenging; it did well up to and including dimension 8, but did not perform very well in higher dimensions. The MSE for the first LPN decreased by a large factor in $d=64$ compared to $d=32$, and again we are unsure why this is the case. Nonetheless, Fig. 5 shows that the learned prior turned out to approximate the true backward viscosity solution given by Eq. 18, as expected from the theory of backward viscosity solutions.
In addition to our numerical experiments, we show below the performance of the regularized inversion method proposed in [fang2024whats] to estimate $J$ from the first LPN [^3] Find $\bm{y}$ such that $\text{prox}_{J}(\bm{y})=\bm{x}$ by minimizing $\psi(\bm{y})+\frac{\alpha}{2}\left\|{\bm{y}}\right\|_{2}^{2}-\langle\bm{x},\bm{y}\rangle)$ w.r.t. $\bm{y}$ with $\alpha\geqslant 0$.. As Fig. 7 shows, the regularized inversion method performed very poorly and we thus do not comment on it further.
@@
Figure 5: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension 44 (left). The cross sections of the prior function learned from our trained second LPN method (right).
Figure 6: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension 88 (left). The cross sections of the prior function learned from our trained second LPN method (right).
Figure 7: The cross sections of the prior function learned from the regularized inversion method proposed in [fang2024whats] Dimension 44 (left) and dimension 88 (right).
@@ 6 Discussion
In this work, we leveraged the theory of viscosity solutions of HJ PDEs and recent work on proximal operators to develop novel deep learning numerical methods for learning the underlying prior of the proximal operator Eq. 2 from data. Our approach built on existing connections between proximal operators and HJ PDEs, namely that proximal operators are obtained from solutions to a class of HJ PDEs, and on the theory for the inverse problem for HJ equations. As discussed in Section 3, while there may be infinitely many priors that can recover Eq. 1, there is a natural choice, obtained by reversing the time in the HJ PDE Eq. 14 and using the value of the proximal operator $(\bm{x},t)\mapsto S(\bm{x},t)$ as initial condition. The resulting backward viscosity solution yields a semiconvex prior $J_{\text{BVS}}$ that can reconstruct the value function $(\bm{x},t)\mapsto S(\bm{x},t)$ underlying proximal operator. We considered the special case where only samples of the proximal operators and its values were available in Section Section 4, and used techniques from max-plus algebra to derive some characterizations and errors property of $J_{\text{BVS}}$ with respect to convex functions approximating it from above. Finally, in Section 5 we presented numerical experiments to learn both the forward solution $\bm{x}\mapsto S(\bm{x},t)$ and the prior $J_{\text{BVS}}$ by training a convex neural network, specifically a learned proximal network from data $\{\bm{x}_{k},S(\bm{x}_{k},t)\}_{k=1}^{K}$. We presented several numerical results that demonstrate the efficiency of our proposed method in high dimensions.
@@
Appendix A Formal calculation of Eq. 23